# Problem-informed Graphical Quantum Generative Learning

### v1

@@ III.1 From Graphical Representation to Variational Ansatz
where $Z_{v}$ is the Pauli-$Z$ operator acting on qubit $v$, $\mathcal{C}_{\mathcal{H}}$ refers to the set of cliques and $\boldsymbol{\beta}$ is the set of parameters. Usually some of the MN cliques overlap in nonzero subsets, thus there will be reoccurring terms. Since all terms commute, we can reparametrize the Hamiltonian such that each term only appears once (and identities are excluded): $H^{\prime}(\boldsymbol{\beta})\rightarrow H(\boldsymbol{\alpha})$.
@@
and implement a model similar to QCIBMs, defined in Sec. II.2.1 and Fig. 3. We call these problem-informed QCBM models quantum circuit Markov random fields (QCMRFs).
@@
Alternatively, one can also limit the locality of the Hamiltonian to get shallower circuits, that in turn can be worse at capturing higher-order correlations. This is ultimately equivalent to considering smaller cliques instead of the maximal clique factorization. The circuit implementation details are discussed in Appendix A.1.
@@ III.2 Benchmark Proposal
Benchmarking generative QML models often relies on generic probability distributions, such as the bars and stripes dataset, or some Hamming weight specific target distribution [11, 42, 49]. Here we describe our proposal for constructing target distributions for these models, where the complexity is tunable in several ways. This construction relies on MNs, where the graph structure can be defined by the user. In general, the “difficulty” of the learning problem is proportional to the clique sizes of the MN. The most general case is a complete graph with a single maximal clique, that corresponds to explicitly encoding the probability to each global state.
Given an undirected graph $\mathcal{G}$, a set of cliques $\mathcal{C}_{\mathcal{H}}$, and a classical generator, we construct a target MN $\mathcal{H}$ through the following steps:
@@
    To each clique $C\in\mathcal{C}_{\mathcal{H}}$, we assign a factor as a set of $2^{|C|}$ numbers obtained by querrying the classical generator $2^{|C|}$ times.
@@
    Next, we calculate the unnormalized measures of each global assignment by multiplying the corresponding element of each factor.
@@
As the size of the graph increases, this procedure becomes highly inefficient, because of the exponential size of the space. In these cases, instead of calculating the target probabilities exactly, steps $2-4$ can be replaced by the use of approximate sampling techniques to sample $\mathcal{H}$ directly [30], e. g., using Gibbs sampling.
@@ III.3.1 Benchmarks against QCIBMs
The results are shown in the top row of Fig. 4. We start with a $3\times 3$ grid, which defines a pairwise MN, meaning, that the corresponding QCMRF model incorporates only $2$-local interactions. Here both models have similar performance, while our QCMRF model reduces the number of trainable parameters from $72$ to $48$.
@@
This series of experiments show, that as we increase the connectivity of the graph and the sizes of its maximal cliques, the distribution becomes harder to learn, as it is reflected in the performance of the problem-agnostic QCIBM. However, the performance of our problem-specific QCMRF model is either unaffected by this change, or it performs even better, as its complexity also increases with the underlying MN. This also proves, that using higher order Hamiltonians can actually help in capturing higher order correlations between the random variables of the MN.
@@
III.3.2 Benchmarks againts BBQCs
Next, we validate our model with BBQCs on non-chordal MNs. In these cases, the undirected graph first has to be triangulated and turned into a directed acyclic graph in order to consider the corresponding BN. Knowing the structure of this BN, one can implement the corresponding BBQC, that is supposed to be able to capture the target distribution exactly. However, this process is very costly: the triangulation of the graph itself is a hard problem and it can introduce a large number of edges, that leads to the introduction of additional variables. Consequently, the corresponding BBQC can have many more trainable parameters and the circuit depth can be significantly larger.
For this comparison, we considered loop graphs, that are easy to triangulate. Since with 00 parameter initialization, the starting probability distribution is different for QCMRF and BBQC circuits, here we started with random values to have a more fair comparison. The number of shots was set to $10^{3}$ for all experiments along with the size of the training set for the MMD-based optimizations. We ran all experiments with $10$ sets of uniformly random factor values and visualized the average performance in Fig. 5.
All these loop-graphs can be used to implement QCMRF circuits, while to define the corresponding BBQC, the triangulation introduces $n-3$ new edges. Due to this fact, the number of trainable parameters along with the depth of the circuit is significantly increased. In all cases, our MN-based model reaches similar performance to BBQCs, while having much lower cost of implementation, which further highlights the usefulness of our model class. For a deeper comparison between the implementation costs of these models, we refer the reader to Appendix B.
@@ III.4 Trainability
A complete graph having maximal clique factorization corresponds to no problem-specific knowledge, since the probability distribution has $12^{n}-$ 1 degrees of freedom. As shown in Fig. 6, according to our numerics, the cost function variance vanishes exponentially in the number of qubits in this case, which also indicates the presence of deterministic barren plateaus [52]. To study sparser (but still dense) graphs, we concentrate first on Erdős-Rényi graphs with $p=0.5$ edge probability, since in this case, the expected size of the largest clique is $\mathcal{O}(\log n)$ with high probability. Our simulations clearly show, that the scaling of the variance is closer to polynomial. This is also true for triangle chain graphs, where we have $n-2$ cliques of size $3$ in total. While polynomial scaling of the cost variance does not prove the empirical absence of barren plateaus, it is a good first indicator of better trainability.
@@
It follows naturally from the our definitions, that QCMRF circuits corresponding to efficient MNs representations, have depth $\mathrm{poly}(n)$.
@@ III.5 Potential for Quantum Advantage
Quantum advantage, in the context of generative learning, can have various flavours: it can show improvement in the precision (wrt a distance metric); it can refer to faster convergence; or even an improvement in the number of training samples needed. In order to formulate all these cases in a single definition, we first need to define what we mean by a class of distributions being efficiently learnable:
@@
An algorithm $\mathcal{A}\in BQP$ is said to have quantum learning advantage, if there exists a class of distributions $P_{n}$ for which there exists $d$, $\epsilon$, $k$ such that $P^{n}$ is $(d,\varepsilon,k,BQP)$-learnable, but not $(d,\varepsilon,k,BPP)$-learnable.
@@
In the following, we concentrate on this second definition and argue, that since our QCMRF model class contains the class of QAOA and IQP circuits, it can also produce distributions that are thought to be classically hard. For this, we assume, that the joint distribution of a target MN is learnable by both an arbitrary classical model and a QCMRF model to a given precision $\varepsilon$ and analyze the complexity of sampling the trained models. Previous works [13, 53] presented similar arguments relying on the results of Refs. [54, 7]. We start by sketching current results regarding the hardness of sampling QAOA circuits, then based on this, we present our conjecture about the possible quantum advantage in sampling for our model.
The Quantum Approximate Optimization Algorithm (QAOA) was proposed in [36] for approximately solving combinatorial optimization problems on a quantum computer. In [7], the authors proved, that efficient classical sampling of the output distribution of the $1$-level QAOA circuit implies the collapse of the polynomial hierarchy to the third level. While the argument was constructed with a $2$-local QAOA circuit, the proof stands for higher order $Z$ interactions as well, all these being diagonal operators. On the other hand, this was shown for multiplicative error only, meaning, that if the target distribution (being the one defined by the QAOA circuit) $P_{QAOA}$ and the classical model distribution $P_{\mathcal{M}}$ satisfy the bound
@@
then the polynomial hierarchy collapses to its third level. In our framework, however, we mostly concentrated on distance as measured in TV, which is equivalent to additive error. Multiplicative error is a stronger constraint, thus simulating up to a bounded additive error is a harder task. These results using multiplicative error were extended in [55] from worst case to average case hardness for $1$-level QAOA circuits. Finally, in [37], the author proved average case hardness with additive error, which, up to our knowledge, is the strongest result in connection with the weak simulation of $1$-level QAOA circuits. Since the class of probability distributions defined by the output of QCMRF circuits contains the $1$-level QAOA circuits, this larger class can also be hard to weakly simulate, i.e., to sample the output probability distribution efficiently classically.
Figure 7: The conjectured classes of probability distributions. We know, that the intersection of BNs and MNs are the class of chordal graphs (see Sec. II.1) and that the class of QCMRF circuits contains the class of QAOA circuits. We conjecture however, that the PGM classes and that of the quantum circuits mentioned also have a nonzero intersection, that contains classically hard cases.
The classical hardness of distributions efficiently captured by MNs and QCMRF circuits makes it reasonable to believe that these two classes have a nonzero intersection that also contains hard cases. This also means, that - provided that we can train a QCMRF model to sufficient precision - we can use the trained model to efficiently sample the distribution of the underlying MN. These facts together with our numerical findings lead to the following conjecture, that is also illustrated in Fig. 7.
@@
The class of QCMRF circuits, that can learn probability distribution efficiently represented by MNs, also contains classically hard cases, yielding a quantum advantage in sampling.
@@ IV Conclusion & Outlook
Our numerical experiments demonstrated that our model, called quantum circuit Markov random field, is capable of capturing higher order correlations between the binary random variables of the corresponding MN. This can significantly improve the performance in the case of higher order target models, while potentially reduce the number of trainable parameters on sparse graphs, compared to problem-agnostic approaches. We further validated our model with basis-enhanced Bayesian quantum circuits on non-chordal MNs, since these Bayesian network-based models are able to express the target distribution exactly. The QCMRF models reached the performance of BBQCs on small loop-graphs with less parameters and significantly shallower circuits. All these experiments were conducted using the KL divergence and MMD loss functions, to demonstrate both exact distribution learning tasks and more practical generative learning based on limited training samples.
A preliminary numerical analysis of trainability was presented, that introduced an important constraint on the sparsity of the potentential MNs of interest. We formulated two definitions of quantum advantage, relevant in the context of generative models, where the first concentrated on a learning advantage, and the second focused on efficiently sampling from the learned distribution. We presented an argument in the second setting, highlighting the potential of our model to offer improvements over classical methods, since it contains the class of QAOA circuits, that are known to be hard to sample from. We also believe that this connection between classical MNs and QAOA-type circuits opens up an interesting direction for further investigations.
While we concentrated on learning a target distribution to high accuracy, this alone is not enough to characterize the performance of generative models. Another important factor is the model’s ability to generalize, rather than memorize the training data [56]. In the context of PGMs, this can be investigated with a training dataset of limited size, assessing the trained model’s ability to generate valid but unseen samples.
We also remark, that while our QCMRF model shows significant improvement compared to problem-agnostic QCIBMs on several small examples, BBQCs show better performance on chordal graphs, where both PGM-based models have the same number of trainable parameters. This means, that for chordal graphs, BN-based models are better at capturing the target distribution, although requiring much deeper circuits.
@@ V Acknowledgements
The authors would like to thank the support of the Hungarian National Research, Development and Innovation Office (NKFIH) through the KDP-2021 and KDP-2023 funding scheme, the Quantum Information National Laboratory of Hungary and the grants TKP-2021-NVA-04 and FK 135220. The authors also acknowledge the computational resources provided by the Wigner Scientific Computational Laboratory (WSCLAB).

### v2

@@ III.1 From Graphical Representation to Variational Ansatz
where $Z_{v}$ is the Pauli-$Z$ operator acting on qubit $v$, $\mathcal{C}_{\mathcal{H}}$ refers to the set of cliques and $\boldsymbol{\beta}$ is the set of parameters. Please note, that the $\boldsymbol{\beta}$ parameters are only used to enable a compact representation, but ultimately do not reflect the number of free parameters in the final model. Usually some of the MN cliques overlap in non-zero subsets, thus there will be reoccurring terms. Since all terms commute, we can reparametrize the Hamiltonian such that each term only appears once with a distinct coefficient (and identities are excluded): $H^{\prime}(\boldsymbol{\beta})\rightarrow H(\boldsymbol{\alpha})$. Also note, that in this new parametrization, each element of $\boldsymbol{\alpha}$ is independent, in the most general case, and $|\boldsymbol{\alpha}|\geq|\boldsymbol{\beta}|$.
@@
and implement a model similar to QCIBMs, defined in Sec. II.2.1 and Fig. 3. We call these problem-informed QCBM models quantum circuit Markov random fields (QCMRFs). For a more detailed model construction process, please refer to Appendix A, where we rely on the representation of probability distributions as diagonal matrices and use the Pauli expansion of the projectors. Furthermore, we also show that the output probability distribution of the QCMRF model can be factorized according to the same MN, that was used in its construction.
@@
Alternatively, one can also limit the locality of the Hamiltonian to get shallower circuits, that in turn can be worse at capturing higher-order correlations. This is ultimately equivalent to considering smaller cliques instead of the maximal clique factorization. The circuit implementation details are discussed in Appendix B.1.
@@ III.2 Benchmark Proposal
Benchmarking generative QML models often relies on generic probability distributions, such as the bars and stripes dataset, or some Hamming weight specific target distribution [11, 43, 50]. Here, we describe our proposal for constructing target distributions for these models, where the complexity is tunable in several ways. This construction relies on MNs, where the graph structure can be defined by the user. In general, the “difficulty” of the learning problem is proportional to the clique sizes of the MN. The most general case is a complete graph with a single maximal clique, that corresponds to explicitly encoding the probability to each global state.
Given an undirected graph $\mathcal{G}$, a set of cliques $\mathcal{C}_{\mathcal{H}}$, and a classical (pseudo)random number generator, we construct a target MN $\mathcal{H}$, as in Def. 2, through the following steps:
@@
    To each clique $C\in\mathcal{C}_{\mathcal{H}}$, we assign a factor as a set of $2^{|C|}$ numbers obtained by querying the classical generator $2^{|C|}$ times.
@@
    Next, we calculate the unnormalized measures of each global assignment by multiplying the corresponding element of each factor table.
@@
As the size of the graph increases, this procedure becomes highly inefficient, because of the exponential size of the space. In these cases, instead of calculating the target probabilities exactly, steps $2-4$ can be replaced by the use of approximate sampling techniques to sample $\mathcal{H}$ directly [31], e.g., by using Markov chain Monte Carlo methods.
@@ III.3.1 Benchmarks against QCIBMs
The results are shown in the top row of Fig. 4. We start with a $3\times 3$ grid, which defines a pairwise MN, meaning, that the corresponding QCMRF model incorporates only $2$-local interactions. Here, both models have similar performance, while our QCMRF model reduces the number of trainable parameters from $72$ to $48$.
@@
This series of experiments show, that as we increase the connectivity of the graph and the sizes of its maximal cliques, the distribution becomes harder to learn, as it is reflected in the performance of the problem-agnostic QCIBM. However, the performance of our problem-specific QCMRF model is either unaffected by this change, or it performs even better, as its complexity also increases with the underlying MN. This also proves, that using higher-order Hamiltonians can actually help in capturing higher-order correlations between the random variables of the MN.
@@
We also showcase the scaling of the performance separation with the number of qubits for a triangle chain graph. These results based on exact simulation (infinite shot limit) are shown in Appendix D, where we also present a minimal numerical comparison to classical methods.
III.3.2 Benchmarks against BBQCs
Next, we validate our model with BBQCs on non-chordal MNs. In these cases, the undirected graph first has to be triangulated and turned into a directed acyclic graph. Knowing the structure of this BN, one can implement the corresponding BBQC, that is able to capture the target distribution exactly. However, this process is very costly: the triangulation of the graph itself is a hard problem, and it can introduce a large number of edges, that leads to the introduction of additional variables. Consequently, the corresponding BBQC can have many more trainable parameters and the circuit depth can be significantly larger.
For this comparison, we considered loop graphs, that are easy to triangulate. Since with 00 parameter initialization, the starting probability distribution is different for QCMRF and BBQC circuits, here we started with random values to have a more fair comparison. The number of shots was set to $10^{3}$ for all experiments, along with the size of the training set for the MMD-based optimizations. We ran all experiments with $10$ sets of uniformly random factor values and visualized the average performance in Fig. 5.
All these loop-graphs can be used to implement QCMRF circuits, while to define the corresponding BBQC, the triangulation introduces $n-3$ new edges. Due to this fact, the number of trainable parameters along with the depth of the circuit is significantly increased. In all cases, our MN-based model reaches similar performance to BBQCs, while having much lower cost of implementation, which further highlights the usefulness of our model class. For a deeper comparison between the implementation costs of these models, we refer the reader to Appendix C.
@@ III.4 Trainability
A complete graph having maximal clique factorization corresponds to no structure, since the probability distribution has $12^{n}-$ 1 degrees of freedom. As shown in Fig. 6, according to our numerics, the cost function variance vanishes exponentially in the number of qubits in this case, which also indicates the presence of deterministic barren plateaus [53]. To study sparser (but still dense) graphs, we concentrate first on Erdős-Rényi graphs with $p=0.5$ edge probability, since in this case, the expected size of the largest clique is $\mathcal{O}(\log n)$ with high probability. Our simulations clearly show, that the variance exhibits much better scaling in this case, than the conclusively exponential decay before. This is also true for triangle chain graphs, where we have $n-2$ cliques of size $3$ in total. While cost concentration and gradient concentration go hand in hand in variational quantum algorithms, polynomial scaling in the cost variance can be induced by a few parameters with good behaviour [53, 54]. Consequently, such a polynomial scaling does not show the empirical absence of barren plateaus, but it is a good first indicator of better trainability.
@@
It follows naturally from our definitions, that QCMRF circuits, corresponding to efficient MN representations, have depth $\mathrm{poly}(n)$.
@@ III.5 Potential for Quantum Advantage
Quantum advantage, in the context of generative learning, can have various flavours: it can show improvement in the precision (with respect to a distance metric); it can refer to faster convergence; or even an improvement in the number of training samples needed. In order to formulate all these cases in a single definition, we first need to define what we mean by a class of distributions being efficiently learnable:
@@
An algorithm $\mathcal{A}\in BQP$ is said to have quantum learning advantage, if there exists a class of distributions $P^{n}$ for which there exists $d$, $\epsilon$, $k$ such that $P^{n}$ is $(d,\varepsilon,k,BQP)$-learnable, but not $(d,\varepsilon,k,BPP)$-learnable, where this concept of learnability is defined in Def. 4.
@@
In the following, we concentrate on this second definition and argue, that since our QCMRF model class contains the class of QAOA and IQP circuits, it can also produce distributions that are thought to be classically hard. For this, we assume, that the joint distribution of a target MN is learnable by both an arbitrary classical model and a QCMRF model to a given precision $\varepsilon$ and analyse the complexity of sampling the trained models. Previous works [13, 55] presented similar arguments, relying on the results of Refs. [56, 7]. We start by sketching current results regarding the hardness of sampling QAOA circuits, then based on this, we present our conjecture about the possible quantum advantage in sampling for our model.
The Quantum Approximate Optimization Algorithm (QAOA) was proposed in Ref. [37] for approximately solving combinatorial optimization problems on a quantum computer. In [7], the authors proved, that efficient classical sampling of the output distribution of the $1$-level QAOA circuit implies the collapse of the polynomial hierarchy to the third level. While the argument was constructed with a $2$-local QAOA circuit, the proof stands for higher-order $Z$ interactions as well, all these being diagonal operators. On the other hand, this was shown for multiplicative error only, meaning, that if the target distribution (being the one defined by the QAOA circuit) $P_{QAOA}$ and the classical model distribution $P_{\mathcal{M}}$ satisfy the bound
@@
then the polynomial hierarchy collapses to its third level. In our framework, however, we mostly concentrated on distance as measured in TV, which is equivalent to additive error. Multiplicative error is a stronger constraint, thus simulating up to a bounded additive error is a harder task. These results using multiplicative error were extended in Ref. [57] from worst case to average case hardness for $1$-level QAOA circuits. Finally, in Ref. [38], the author proved average case hardness with additive error, which, up to our knowledge, is the strongest result in connection with the weak simulation of $1$-level QAOA circuits. Since the class of probability distributions defined by the output of QCMRF circuits contains the $1$-level QAOA circuits, this larger class can also be hard to weakly simulate, i.e., to sample the output probability distribution efficiently classically. This also means, that - provided that we can train a QCMRF model to sufficient precision - we can use the trained model to efficiently sample the distribution of the underlying MN. These facts together with our numerical findings lead to the following conjecture.
@@
The class of QCMRF circuits, that can learn probability distributions efficiently represented by MNs, also contains classically hard cases, yielding a quantum advantage in sampling.
@@ IV Conclusion & Outlook
Our numerical experiments demonstrated that our model, called quantum circuit Markov random field, is capable of capturing higher-order correlations between the binary random variables of the corresponding MN. This can significantly improve the performance in the case of higher-order target models, while potentially reduce the number of trainable parameters on sparse graphs, compared to problem-agnostic approaches. We further validated our model with basis-enhanced Bayesian quantum circuits on non-chordal MNs, since these BN-based models are able to express the target distribution exactly. The QCMRF models reached the performance of BBQCs on small loop-graphs with fewer parameters and significantly shallower circuits. All these experiments were conducted using the KL divergence and MMD loss functions, to demonstrate both exact distribution learning tasks and more practical generative learning based on limited training samples.
A preliminary numerical analysis of trainability was presented, that introduced an important constraint on the sparsity of the potential MNs of interest. We formulated two definitions of quantum advantage, relevant in the context of generative models, where the first concentrated on a learning advantage (Def. 5), and the second focused on efficiently sampling from the learned distribution (Def. 6). We presented an argument in the second setting, highlighting the potential of our model to offer improvements over classical methods, since it contains the class of QAOA circuits, that are known to be hard to sample classically. We also believe that this connection between classical MNs and QAOA-type circuits opens up an interesting direction for further investigations.
While we concentrated on learning a target distribution to high accuracy, this alone is not enough to characterize the performance of generative models. Another important factor is the model’s ability to generalize, rather than memorize the training data [58]. In the context of PGMs, this can be investigated with a training dataset of limited size, assessing the trained model’s ability to generate valid but unseen samples. We also remark, that while our QCMRF model shows significant improvement compared to problem-agnostic QCIBMs on several small examples, BBQCs show better performance on chordal graphs, where both PGM-based models have the same number of trainable parameters. This means, that for chordal graphs, BN-based models are better at capturing the target distribution, although requiring much deeper circuits. Finally, what can be seen as a major setback of our framework is that it assumes knowledge of the graph structure along with its cliques. While this seems like a major obstacle, this knowledge is readily available in various use cases. Furthermore, this generative QML approach can be used together or as a subroutine of classical structure learning algorithms [59].
@@ Acknowledgements
The authors would like to thank the support of the Hungarian National Research, Development and Innovation Office (NKFIH) through the KDP-2021 and KDP-2023 funding scheme, the Quantum Information National Laboratory of Hungary and the grants TKP-2021-NVA-04 and FK 135220. Z.Z. was partially supported by the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100 and the QuantERA II project HQCC-101017733. The authors also acknowledge the computational resources provided by the Wigner Scientific Computational Laboratory (WSCLAB).
Appendix A Ansatz design
Here we give a more detailed description of our QCMRF Ansatz design process. Let us consider a single clique $C$, with factor $\phi_{C}$, and embed the corresponding marginal probability distribution in the diagonal of a $2^{|C|}\times 2^{|C|}$ matrix as
$H_{C}$ $=\mathrm{diag}(P_{\phi_{C}})$ (11)
$=\frac{1}{Z_{\phi_{C}}}\sum_{x<2^{|C|}}\phi_{C}(x)\outerproduct{x}{x}$ |
$=\frac{1}{Z_{\phi_{C}}}\sum_{x<2^{|C|}}\phi_{C}(x)\bigotimes_{i}\frac{1}{2}(\mathds{1}_{i}+(-1)^{x_{i}}Z_{i}).$
Here $Z_{\phi_{C}}$ denotes the normalizing constant of the marginal distribution, $|C|$ denotes the size of clique $C$, and in the last line, we used the Pauli expansion of the projectors.
We can reparametrize this linear combination of Pauli strings as
$\begin{split}H_{C}(\boldsymbol{\alpha})&=\alpha_{0}\bigotimes_{i}\mathds{1}_{i}+\alpha_{1}Z_{0}+\alpha_{2}Z_{1}+\dots\\ &+\alpha_{2^{|C|}-1}\bigotimes_{i}Z_{i}.\end{split}$ (12)
After discarding the identity term, we consider the unitary $U_{C}(\boldsymbol{\alpha})=\exp(-iH_{C}(\boldsymbol{\alpha}))$. At this point, it is easy to see, that if this operator acts on the initial state as $U_{C}(\boldsymbol{\alpha})\ket{+}^{\otimes n}$, then the marginal probability distribution is encoded into the phases of each computational basis state. We can also see, that there are $12^{|C|}-$ 1 parameters, which is exactly how many degrees of freedom a probability distribution over $|C|$ binary random variables has.
Going back to the full probability distribution $P_{\Phi}$, that factorises according to a MN as in Def. 2, we can embed this into the diagonal of a $2^{n}\times 2^{n}$ matrix of the form
$H=\mathrm{diag}(P_{\Phi})$ $=\frac{1}{Z_{\Phi}}\prod_{C\in\mathcal{C}}\left[\sum_{x<2^{|C|}}\phi_{C}(x)\bigotimes_{i\in C}\frac{1}{2}(\mathds{1}_{i}+(-1)^{x_{C(i)}}Z_{i})\bigotimes_{i\notin C}\mathds{1}_{i}\right]$ (13)
$=\frac{1}{Z_{\Phi}}\sum_{x<2^{n}}\left[\prod_{C\in\mathcal{C}}\phi_{C}(x_{C})\bigotimes_{i\in C}\frac{1}{2}(\mathds{1}_{i}+(-1)^{x_{i}}Z_{i})\bigotimes_{i\notin C}\mathds{1}_{i}\right].$
The first two lines show equivalent definitions, where $C(i)$ denotes the index of node $i$ in clique $C$ and $x_{C}$ denotes the bits in the binary form of $x$, that correspond to clique $C$.
According to the previous argument, after reparametrizing this Hamiltonian as before, the unitary $U_{Z}(\boldsymbol{\alpha})=\exp(-iH(\boldsymbol{\alpha}))$ is expressive enough to encode the marginal probability distributions over each clique. However, since the factors are not normalized in general, the joint distribution is not necessarily the same as the factor product of the marginal ones. For this reason, instead of extracting the phases of the state $U\ket{+}^{\otimes n}$, we perform a measurement on each qubit in any local basis, i.e., we apply general one-qubit gates before measurement, where these gates have independent trainable parameters. We expect, that this expressivity enhancement provides enough flexibility to compensate for the mismatch coming from the different scale of the marginal distributions.
We can also take the opposite route and analyse the output probability distribution of this quantum circuit model. We want to analyze the structure of the MN, that factorizes this probability distribution. First, we can write the marginal probability of each random variable in the output distribution as
$P(X_{i})=-\frac{\langle Z_{i}\rangle-1}{2},$ (14)
where $\langle Z_{i}\rangle$ denotes the expectation value of the Pauli-Z operator on qubit $i$. Thus, it suffices to examine the backward light-cone of observable $Z_{i}$ for all $i$. Due to the structure of the QCMRF model, the entangling gates in the $U_{Z}$ operator can be ordered in such a way, that the light-cone of the local observable $Z_{i}$ only contains the qubits, it is directly connected to, through multi-qubit gates. This means, that the probability in Eq.(14) can be written as a product of probabilities and conditional probabilities involving only the neighbours of node $i$. This is equivalent to the Markov property, stating that random variable $X_{i}$ is conditionally independent of all other variables given its neighbours.
Since in $U_{Z}$, all operator terms commute, each ordering of the gates (generated by the tensor product of Pauli-Z operators) produces the same quantum state leading to the same output probability distribution. Therefore, the Markov property applies to all output random variables simultaneously. Due to the Hammersley-Clifford theorem for strictly-positive probability distributions [60], the presence of this Markov property also means, that the joint probability distribution factorizes according to the corresponding MN. Consequently, the output probability distribution of QCMRF circuits factorizes according to the same MN structure, assuming positive probabilities.
Appendix B Implementation details